# GA-Drive: Geometry-Appearance Decoupled Modeling for Free-viewpoint Driving Scene Generation

### v1

@@ 3 Method
Fig.2 shows the overview of our framework. The geometry module is presented in Sec. 3.1. We then describe the Geometry-Guided Pseudo-View Synthesis in Sec. 3.2 and the Video Diffusion Model in Sec. 3.3. Finally, the Pseudo-View Simulation Pipeline for training is detailed in Sec. 3.4.
@@ 3.2 Geometry-Guided Pseudo-View Synthesis
Given a 4D reconstruction of the driving scene, we cast a ray from each pixel into 3D space based on the pixel’s depth, which is derived from the 4D reconstruction. At a particular timestamp, assume we have access to $N$ recorded images , along with their corresponding camera poses and depth maps rendered from the Gaussian field. Given a target camera pose along a novel trajectory, our objective is to synthesize a pseudo-view by leveraging these $N$ source images.
Problem Setup  Given $N$ source views $\{(\mathbf{I}_{i},\mathbf{K}_{i},\mathbf{T}_{i},\mathbf{D}_{i})\}_{i=1}^{N}$, where $I_{i}$ is the known RGB image, $\mathbf{K}_{i}\in\mathbf{R}^{3\times 3}$ is the camera intrinsic matrix, $\mathbf{T}_{i}\in SE(3)$ is the camera-to-world transformation, $\mathbf{D}_{i}\in\mathbf{R}^{H\times W}$ is the depth map rendered from the gaussian field, and a target camera defined by $(\mathbf{K}_{t},\mathbf{T}_{t},\mathbf{D}_{t})$. Our goal is to synthesize a pseudo-view $I_{t}^{\text{pseudo}}$ that approximates the appearance of the scene from the target viewpoint.
@@ 3.3 Segment-wise Video Diffusion Model
Our video diffusion model builds upon the I2V architecture from CogVideoX [24]. To incorporate conditioning frames in the video diffusion model, we encode them using the same 3D variational autoencoder (VAE) as CogVideoX. The resulting latent features are then concatenated with Gaussian noise along the channel dimension. A segment of conditioning frames is represented as a tensor of shape $T\times H\times W\times 3$, where $T$ is the number of frames, $H$ and $W$ are the height and width, and $3$ denotes RGB channels. The 3D VAE encodes this to a latent tensor of size $\frac{T}{q}\times\frac{H}{p}\times\frac{W}{p}\times 16$, with $q$ and $p$ as temporal and spatial downsampling factors and $16$ latent channels. We sample a gaussian noise tensor of the same shape, concatenate it with the encoded features along the channel dimension, yielding a combined tensor of shape $\frac{T}{q}\times\frac{H}{p}\times\frac{W}{p}\times 32$, which is input to the diffusion model’s denoising network. Model fine-tuning follows the DDPM framework [9], as in CogVideoX. In our setup, $T=16$, $H=640$, $W=960$, and the text prompt is fixed as “A realistic autonomous driving scene.”

### v2

@@ 3 Method
Fig.2 shows the overview of our framework. The geometry module is presented in Sec. 3.1. We then describe the Geometry-Guided Pseudo-View Synthesis in Sec. 3.2 and the Video Diffusion Model in Sec. $3.$ Finally, the Pseudo-View Simulation Pipeline for training is detailed in Sec. 3.4.
@@ 3.2 Geometry-Guided Pseudo-View Synthesis
Given a 4D reconstruction of the driving scene, we cast a ray from each pixel into 3D space based on the pixel’s depth, which is derived from the 4D reconstruction. At a particular timestamp, assume we have access to NN recorded images , along with their corresponding camera poses and depth maps rendered from the Gaussian field. Given a target camera pose along a novel trajectory, our objective is to synthesize a pseudo-view by leveraging these NN source images.
Problem Setup  Given NN source views $\{(\mathbf{I}_{i},\mathbf{K}_{i},\mathbf{T}_{i},\mathbf{D}_{i})\}_{i=1}^{N}$, where $I_{i}$ is the known RGB image, $\mathbf{K}_{i}\in\mathbf{R}^{3\times 3}$ is the camera intrinsic matrix, $\mathbf{T}_{i}\in SE(3)$ is the camera-to-world transformation, $\mathbf{D}_{i}\in\mathbf{R}^{H\times W}$ is the depth map rendered from the gaussian field, and a target camera defined by $(\mathbf{K}_{t},\mathbf{T}_{t},\mathbf{D}_{t})$. Our goal is to synthesize a pseudo-view $I_{t}^{\text{pseudo}}$ that approximates the appearance of the scene from the target viewpoint.
@@ 3.3 Segment-wise Video Diffusion Model
Our video diffusion model builds upon the I2V architecture from CogVideoX [24]. To incorporate conditioning frames in the video diffusion model, we encode them using the same 3D variational autoencoder (VAE) as CogVideoX. The resulting latent features are then concatenated with Gaussian noise along the channel dimension. A segment of conditioning frames is represented as a tensor of shape $T\times H\times W\times 3$, where $T$ is the number of frames, $H$ and $W$ are the height and width, and 33 denotes RGB channels. The 3D VAE encodes this to a latent tensor of size $\frac{T}{q}\times\frac{H}{p}\times\frac{W}{p}\times 16$, with $q$ and $p$ as temporal and spatial downsampling factors and 1616 latent channels. We sample a gaussian noise tensor of the same shape, concatenate it with the encoded features along the channel dimension, yielding a combined tensor of shape $\frac{T}{q}\times\frac{H}{p}\times\frac{W}{p}\times 32$, which is input to the diffusion model’s denoising network. Model fine-tuning follows the DDPM framework [9], as in CogVideoX. In our setup, $T=16$, $H=640$, $W=960$, and the text prompt is fixed as “A realistic autonomous driving scene.”